OpenStudy (anonymous):

S=-4s+30 what is the solution to the equation? 1.s=-10 2.s=6 3.s=-6

OpenStudy (angela210793):

5s=30 s=6

OpenStudy (anonymous):

? 2 variables?

OpenStudy (anonymous):

S=-4s+30 +4s +4s 5s = 30 5s/5 = 30/5 S= 6

OpenStudy (anonymous):

if both side s are same then s=6